# Deep Insights into Cognitive Decline: A Survey of Leveraging Non-Intrusive Modalities with Deep Learning Techniques

### v1

@@ V-B1 Recurrent Neural Networks
As previously stated, the ADDReSS dataset consists of audio recordings in which subjects describe The Cookie Theft Picture (see Figure 6). The work in [195] utilizes this image in their approach, where the picture is segmented into distinct areas. For instance, the mother and the child represent separate areas. Additionally, a cluster of synonymous words is proposed for each region (e.g., synonyms for “mother” include “lady” or “woman”). These sets of synonyms, combined with descriptions of the image, are used to construct a matrix that counts co-occurrences in the descriptions. This matrix serves as the input embedding for a subsequent LSTM model, achieving an accuracy of up to 83.6%. The work proposed by Y. Pan et al. [197] combines various methodologies, including LSTMs, an attention layer, and GRU. This method, which surpasses the performance of a single LSTM used for predicting Alzheimer’s disease, achieves an F1-Score of 74.37 on the DementiaBank Pitt Corpus.
@@ V-B2 Transformers
Regarding the works that use the textual modality on the ADReSS dataset, all but one employ Transformer architectures. The work in [87] compares the performance of LongFormer [218] and the base and large versions of BERT. In this case, the best performance was achieved by the LongFormer, with an accuracy of up to 82.02% on the dataset. Moreover, the employment of the large version of BERT also resulted in an improvement over the base version. In the work proposed by I. Loukas et al. [190], the authors experiment with several pre-trained Transformers, including BERT, BioBERT [219], BioClinicalBERT [220], RoBERTa, ConvBERT [221], ALBERT, and XLNet. The most relevant results were obtained using BERT, achieving up to 87.5% accuracy and a 93.33% F1-Score. This work proposed a multiclass model to predict the severity of dementia based on the MMSE Score, but it did not improve the classification approach. The work proposed in [82] employs a BERT model in combination with a SVM, obtaining an accuracy of 85% on ADReSS dataset.
@@
In [61], three different Transformers, RoBERTa, BERT, and AlBERT, were tested on the B-SHARP dataset for the detection of MCI. The RoBERTa model yielded the best individual results. However, combining the embeddings from all three Transformers resulted in even better performance, achieving an accuracy of up to 74.1%. The work proposed in [94] employs the combination of different Transformer embeddings for the text modality, specifically for the detection of MCI in the Taukadial challenge. In this study, the embeddings obtained from a BERT model were integrated with similarity embeddings derived from a comparison between the subjects’ descriptions and the actual descriptions of the images in question. This approach achieves an accuracy of 74.82, an F1-Score of 79.43, and an RMSE score of 3.14. The study by [95] employs BERT embeddings combined with a dense layer to predict MCI using the I-CONECT dataset, achieving an accuracy of 60.75% and an F1-Score of 59.67%. This approach demonstrates an improvement over the results obtained using the audio modality.
@@ V-C Vision modalities
While most research on non-intrusive modalities focuses on audio and text, other relevant modalities can also offer valuable insights for detecting these disorders. Among these are image and video modalities, which together form our proposed visual modalities. The most relevant works within these modalities are shown in Table IV. This table is structured by modality and further categorized by disease and dataset. It presents the year of publication, the methodologies used, and performance metrics such as accuracy and F1-Score. In this case, the RMSE is not reported, as none of the included works provide this information. The best results for each dataset are highlighted in bold. The methodologies are detailed in the following subsections according to the modality applied. These modalities are image and video.
@@
Alzheimer & MCI Private dataset [96] 2022 CNN 72.21 -
@@ V-C2 Video
Another relevant modality for these cognitive estimation tasks is video. Videos provide valuable information from subjects, such as their facial expressions and body poses during different actions, which offer insights into their movements. For instance, the study proposed by J. Sun et al. [249] employs this modality to predict cognitive decline, specifically MCI, using the I-CONNECT dataset. In this work, the authors analyzed the facial expressions and movements of subjects while interacting with others on video recordings. Additionally, features such as lip and eye movements, which play a critical role in both verbal and non-verbal communication, are essential for detecting cognitive impairment [257, 258, 259, 260]. The videos were processed using a video Transformer called MC-ViVit, based on the ViVit transformer for video [117]. With this implementation, the authors achieved an accuracy of 90.63% and an F1-Score of 93.03%. In the same dataset, the approach proposed in [95] utilizes the video modality, specifically raw frames, which are processed using a video Transformer, namely the Video Swin Transformer [261]. This architecture extracts spatio-temporal information from the frames throughout the video. The extracted features are then combined with a dense layer to perform the final classification of the subject’s cognitive state, achieving an accuracy of up to 69.75% and an F1-Score of 66.45%. This modality proves to be the most relevant, as it delivers the best performance when compared to other modalities.
@@
In the work by L. Xing et al. [250], authors propose using video features with their dataset, which contains recordings of individuals using sign language, both those with MCI and healthy individuals. The videos were preprocessed to extract various video features, including facial landmarks and poses. This information was used to train different CNN models, specifically VGG16 and ResNet50. Experiments demonstrated that the VGG16 model achieved the best performance, with an accuracy of 87.88%. Similarly, in the work proposed by D. Abhijit et al. [252] various video features were obtained, including facial action units, gaze, pose, and emotional features. This study focused on their dataset to predict apathy. Although apathy is not a disease, it is common in other conditions and typically affects people’s cognitive abilities. Using the extracted information and a recurrent GRU model, the authors achieved an accuracy of 95.34% and an F1-Score of 94.5%.
@@
Parkinson’s disease has various symptoms, including Freezing of Gait (FoG). FoG is a symptom where patients feel stuck while walking and experience a cessation of movement, even when they intend to keep walking [262, 263]. FoG becomes more frequent over time as the disease progresses, significantly affecting daily life and quality of life [264, 265]. Therefore, detecting this symptom is crucial for identifying Parkinson’s disease and can be achieved through deep learning video analysis [266, 267]. In the work in [251], the authors suggest using pose information from videos of people walking to predict symptoms and the disease in their dataset. This pose information is used to train a recurrent GRU model, which achieves an accuracy of 82.5%.
@@
Eye movement tracking is a technique used to monitor where individuals focus their gaze, providing valuable insights into their visual attention. This data can be obtained through video analysis, similar to how textual data can be derived from the transcription of spoken conversations. In the study proposed in [96], this method has been applied to their dataset to predict Alzheimer’s and MCI. Research has demonstrated a connection between impaired eye movements and cognitive disorders, as these conditions can affect subjects’ attention [268, 269, 270, 271]. In this study, subjects were asked to describe an image that had been segmented into different regions of interest. By analyzing eye movement data, the researchers could determine which areas the subjects focused on while describing the image. This information has been used to train a CNN, specifically the VGG13 model, achieving an accuracy of 72.21%.
@@ VI Multimodality
[77] 2021 CMA A & T 77.2 77.6 -
ADReSS [277] 2022 Joint A & T ? 90.00 89.94 -
[88] 2023 CMA A & T 88.33 88.69 -
[278] 2022 Joint A & T ? 86.25 85.48 -
[82] 2022 Late A & T 86.00 - -
[85] 2020 Late A & T 85.20 85.40 4.65
[87] 2021 Joint A & T 82.92 - -
@@
DementiaBank Pitt Corpus [91] 2023 Joint A & T 86.65 - -
@@
Regarding the pre-existing multimodal models, there are many examples relying on the Transformer architecture, including VideoBERT [282] or VisualBERT [283], among others. The work proposed in [81] experimented with multimodal CLIP [279] and ViLBERT [284]. Additionally, experiments have been carried out on the fusion of ResNet and BERT features obtained from unimodal approaches. These experiments were performed on the ADReSSo dataset to predict Alzheimer’s disease. ViLBERT achieved the best results, with an accuracy and F1-Score of up to 83.7%, outperforming the fusion of single modality approaches.
@@
The other methodology to combine different modalities is the use of fusion strategies. These fusions can be implemented in diverse ways and for various purposes, primarily including early, cross-modal, late, and joint fusion strategies. The goal of these fusions is to integrate different modalities into a single representation of the data.
@@
The early fusion strategy involves combining different sources of information before they are processed by the model [285, 286]. These fusion strategies are also applied in pre-existing multimodal Transformers. One example is the previously mentioned VideoBERT, a Transformer model that utilizes early concatenation to encode the full multimodal context [287]. In [84], an early fusion strategy was used to combine audio and text representations, achieving 81.25% accuracy on the ADReSS dataset and outperforming single-modality approaches.
@@
Cross-modal attention allows interactions across different modalities within a single model [288, 289]. For instance, a cross-modal Transformer encoder combines the key and value vectors of one modality with the query of another [273, 290]. This method, known as cross-attention [291], was first used in the VilBERT model [284]. In [96], the authors introduce a cross-modal attention Transformer that combines eye movement from video and audio features. This approach achieved an accuracy of 84.26% on their dataset, outperforming single modalitiy approaches.
@@
The late fusion strategy involves combining the results of each unimodal model after processing them separately [293]. In [85], the authors propose using majority voting for classification tasks and a weighted average of the results for regression tasks on the ADReSS dataset. This methodology achieved an accuracy of 85.2%, 85.4% F1-Score, and a RMSE score of 4.65. Similarly, the work proposed in [82] also employs majority voting for their unimodal approaches. In this study, this fusion strategy yielded 86% accuracy on the same dataset, outperforming the unimodal approaches.
A joint or hybrid fusion strategy involves combining features obtained from each modality for later use in a final model, but only after being processed individually by each modality model [294]. This combination is the most common approach in the selected works for this study.

### v2

@@ V-B1 Recurrent Neural Networks
As previously stated, the ADDReSS dataset consists of audio recordings in which subjects describe The Cookie Theft Picture (see Figure 6). The work in [194] utilizes this image in their approach, where the picture is segmented into distinct areas. For instance, the mother and the child represent separate areas. Additionally, a cluster of synonymous words is proposed for each region (e.g., synonyms for “mother” include “lady” or “woman”). These sets of synonyms, combined with descriptions of the image, are used to construct a matrix that counts co-occurrences in the descriptions. This matrix serves as the input embedding for a subsequent LSTM model, achieving an accuracy of up to 83.6%. The work proposed by Y. Pan et al. [196] combines various methodologies, including LSTMs, an attention layer, and GRU. This method, which surpasses the performance of a single LSTM used for the detection of Alzheimer’s disease, achieves an F1-Score of 74.37 on the DementiaBank Pitt Corpus.
@@ V-B2 Transformers
Regarding the works that use the textual modality on the ADReSS dataset, all but one employ Transformer architectures. The work in [86] compares the performance of LongFormer [217] and the base and large versions of BERT. In this case, the best performance was achieved by the LongFormer, with an accuracy of up to 82.02% on the dataset. Moreover, the employment of the large version of BERT also resulted in an improvement over the base version. In the work proposed by I. Loukas et al. [189], the authors experiment with several pre-trained Transformers, including BERT, BioBERT [218], BioClinicalBERT [219], RoBERTa, ConvBERT [220], ALBERT, and XLNet. The most relevant results were obtained using BERT, achieving up to 87.5% accuracy and a 93.33% F1-Score. This work proposed a multiclass model to detect the severity of dementia based on the MMSE Score, but it did not improve the classification approach. The work proposed in [81] employs a BERT model in combination with a SVM, obtaining an accuracy of 85% on ADReSS dataset.
@@
In [60], three different Transformers, RoBERTa, BERT, and AlBERT, were tested on the B-SHARP dataset for the detection of MCI. The RoBERTa model yielded the best individual results. However, combining the embeddings from all three Transformers resulted in even better performance, achieving an accuracy of up to 74.1%. The work proposed in [93] employs the combination of different Transformer embeddings for the text modality, specifically for the detection of MCI in the Taukadial challenge. In this study, the embeddings obtained from a BERT model were integrated with similarity embeddings derived from a comparison between the subjects’ descriptions and the actual descriptions of the images in question. This approach achieves an accuracy of 74.82, an F1-Score of 79.43, and an RMSE score of 3.14. The study by [94] employs BERT embeddings combined with a dense layer to detect MCI using the I-CONECT dataset, achieving an accuracy of 60.75% and an F1-Score of 59.67%. This approach demonstrates an improvement over the results obtained using the audio modality.
@@ V-C Vision modalities
While most research on non-intrusive modalities focuses on audio and text, other relevant modalities can also offer valuable insights for detecting these disorders. Among these are image and video modalities, which together form our proposed visual modalities. The most relevant works within these modalities are shown in Table IV. This table is structured by modality and further categorized by disease and dataset. It presents the year of publication, the methodologies used, and performance metrics such as accuracy and F1-Score. The best results for each dataset are highlighted in bold. The methodologies are detailed in the following subsections according to the modality applied. These modalities are image and video. In this case, the RMSE is not reported, as none of the included works provide this information. This absence is partly due to the nature of the datasets and tasks addressed in these studies. Cognitive decline is not a binary state but rather a continuous process that varies in severity. As such, regression tasks based on cognitive assessment scores, such as those from the MMSE or MoCA, are highly relevant, as they enable the estimation of both the degree of decline and the overall cognitive functioning.
Public datasets provided by TalkBank, such as DementiaBank or ADReSS, provide these continuous metrics, making them suitable for such analyses. In contrast, most vision-based studies rely on private datasets due to privacy concerns and limited availability. Additionally, the I-Connect dataset does not include cognitive assessment scores, further limiting its applicability to regression tasks. For these reasons, the works presented in this section do not report RMSE or any regression-based performance metrics.
@@
Alzheimer & MCI
Private dataset [95] 2022 CNN 72.21 -
@@ V-C1 Image
Building on that initial image-based study, several subsequent works by Cilia et al. investigate handwriting in greater depth. First, in [256], authors introduce the DARWIN dataset, the largest public online handwriting corpus for Alzheimer’s detection, and rigorously benchmark task-specific features across multiple classifiers to provide solid baseline results for future research. Second, authors introduce a machine-learning framework that extracts 26 static and dynamic features from individual handwriting strokes, including in-air movements. Using the DARWIN dataset across nine handwriting tasks, they demonstrate that stroke-based analysis performs as well as, or better than, traditional task-level features for predicting Alzheimer’s disease. This insight highlights individual strokes as an informative unit for early cognitive screening [257]. Third, authors extend stroke-level handwriting analysis to 34 tasks and show that combining discriminative feature selection with new ranking-based and stacking ensembles plus SHAP interpretability outperforms aggregated methods for Alzheimer’s detection [258]. Collectively, these studies confirm that handwriting, captured as static images or high-resolution digital traces, offers a non-intrusive, information-rich modality that effectively complements audio and text-based assessments for early cognitive decline. In particular, handwriting analysis can reveal subtle motor planning and visuospatial deficits that may not surface in purely linguistic evaluations.
@@ V-C2 Video
Another relevant modality for these cognitive estimation tasks is video. Videos provide valuable information from subjects, such as their facial expressions and body poses during different actions, which offer insights into their movements. In contrast to image-based approaches that rely on static frames, video-based methods can leverage temporal information to provide richer insights. This temporal dimension allows models to extract information not only from individual visual representations but also from the changes and dynamics across consecutive frames.
For example, video-based analysis is particularly valuable when studying facial cues, as it captures how facial expressions evolve over time. This includes observing how expressions shift in response to spoken discourse, revealing whether certain words or emotional content elicit noticeable facial reactions. In the context of cognitive impairment, such analysis can be especially informative, for instance, by detecting changes in facial expressions as individuals struggle to find the right words. It also extends to microexpressions, subtle facial movements that often occur unconsciously during speech. Furthermore, in cognitive decline estimation tasks, many of which involve picture description, the ability to track whether a person’s gaze or expressions correspond to specific elements of the image being described adds another layer of relevant information. This alignment between visual focus and verbal output can offer valuable cues about cognitive processing.
@@
Regarding cognitive decline estimation, the study proposed by J. Sun et al. [248] employs video modality to detect MCI, using the I-CONNECT dataset. In this work, the authors analyzed the facial expressions and movements of subjects while interacting with others on video recordings. Additionally, features such as lip and eye movements, which play a critical role in both verbal and non-verbal communication, are essential for detecting cognitive impairment [259, 260, 261, 262]. The videos were processed using a video Transformer called MC-ViVit, based on the ViVit transformer for video [116]. With this implementation, the authors achieved an accuracy of 90.63% and an F1-Score of 93.03%. In the same dataset, the approach proposed in [94] utilizes the video modality, specifically raw frames, which are processed using a video Transformer, namely the Video Swin Transformer [263]. This architecture extracts spatio-temporal information from the frames throughout the video. The extracted features are then combined with a dense layer to perform the final detection of the subject’s cognitive state, achieving an accuracy of up to 69.75% and an F1-Score of 66.45%. This modality proves to be the most relevant, as it delivers the best performance when compared to other modalities.
@@
In the work by L. Xing et al. [249], authors propose using video features with their dataset, which contains recordings of individuals using sign language, both those with MCI and healthy individuals. The videos were preprocessed to extract various video features, including facial landmarks and poses. This information was used to train different CNN models, specifically VGG16 and ResNet50. Experiments demonstrated that the VGG16 model achieved the best performance, with an accuracy of 87.88%. Similarly, in the work proposed by D. Abhijit et al. [251] various video features were obtained, including facial action units, gaze, pose, and emotional features. This study focused on their dataset to detect apathy. Although apathy is not a disease, it is common in other conditions and typically affects people’s cognitive abilities. Using the extracted information and a recurrent GRU model, the authors achieved an accuracy of 95.34% and an F1-Score of 94.5%.
@@
Parkinson’s disease has various symptoms, including Freezing of Gait (FoG). FoG is a symptom where patients feel stuck while walking and experience a cessation of movement, even when they intend to keep walking [264, 265]. FoG becomes more frequent over time as the disease progresses, significantly affecting daily life and quality of life [266, 267]. Therefore, detecting this symptom is crucial for identifying Parkinson’s disease and can be achieved through deep learning video analysis [268, 269]. In the work in [250], the authors suggest using pose information from videos of people walking to detect symptoms and the disease in their dataset. This pose information is used to train a recurrent GRU model, which achieves an accuracy of 82.5%.
@@
Eye movement tracking is a technique used to monitor where individuals focus their gaze, providing valuable insights into their visual attention. This data can be obtained through video analysis, similar to how textual data can be derived from the transcription of spoken conversations. In the study proposed in [95], this method has been applied to their dataset to detect Alzheimer’s and MCI. Research has demonstrated a connection between impaired eye movements and cognitive disorders, as these conditions can affect subjects’ attention [270, 271, 272, 273]. In this study, subjects were asked to describe an image that had been segmented into different regions of interest. By analyzing eye movement data, the researchers could determine which areas the subjects focused on while describing the image. This information has been used to train a CNN, specifically the VGG13 model, achieving an accuracy of 72.21%.
@@ VI Multimodality
[76] 2021
CMA
A & T 77.2 77.6 -
ADReSS [279] 2022
Joint
A & T ? 90.00 89.94 -
[87] 2023
CMA
A & T 88.33 88.69 -
[280] 2022
Joint
A & T ? 86.25 85.48 -
[81] 2022
Late
A & T 86.00 - -
[84] 2020
Late
A & T 85.20 85.40 4.65
[86] 2021
Joint
A & T 82.92 - -
@@
DementiaBank
Pitt Corpus
[90] 2023 Joint A & T 86.65 - -
@@
Regarding the pre-existing multimodal models, there are many examples relying on the Transformer architecture, including VideoBERT [284] or VisualBERT [285], among others. The work proposed in [80] experimented with multimodal CLIP [281] and ViLBERT [286]. Additionally, experiments have been carried out on the fusion of ResNet and BERT features obtained from unimodal approaches. These experiments were performed on the ADReSSo dataset to detect Alzheimer’s disease. ViLBERT achieved the best results, with an accuracy and F1-Score of up to 83.7%, outperforming the fusion of single modality approaches.
@@
The other methodology to combine different modalities is the use of fusion strategies. These fusions can be implemented in diverse ways and for various purposes, primarily including early, cross-modal, late, and joint fusion strategies. A visual representation of these fusions is presented in Figure 7. The goal of these fusions is to integrate different modalities into a single representation of the data.
Figure 7: Illustration of modality fusion architectures evaluated in this study: (a) Early Fusion, (b) Late Fusion, (c) Joint Fusion, (d) Cross-Modal Attention Fusion. For simplicity, “M” denotes a Deep Learning model, “Q”, “K”, and “V” denote the query, key, and value, respectively, in the transformer attention mechanism. Blue blocks represent one input modality, green blocks represent the other modality, and purple blocks correspond to fused multimodal representations. Better viewed in color.
@@
The early fusion strategy involves combining different sources of information before they are processed by the model [287, 288]. Early fusion enables the model to capture cross-modal relationships from the outset because the modalities are fused before processing. Several implementations are typical. One is concatenation, where the individual modality representations are concatenated and fed to a shared multimodal module, thereby preserving the entire information from each modality [289]. Another option is to apply simple arithmetic operations, such as element-wise summation or averaging, to obtain a single composite vector. While computationally efficient, these operations blend the modalities and can obscure or distort modality-specific information [290]. This fusion strategy is also applied in pre-existing multimodal Transformers. One example is the previously mentioned VideoBERT, a Transformer model that utilizes early concatenation to encode the full multimodal context [291]. In [83], an early fusion strategy was used to combine audio and text representations, achieving 81.25% accuracy on the ADReSS dataset and outperforming single-modality approaches.
@@
Cross-modal attention allows interactions across different modalities within a single model [292, 293]. Through attention mechanisms, such as those employed in transformers, the model learns to assign varying levels of importance to elements from one modality when interpreting another. In essence, it allows the representation of one modality (e.g., text) to inform and influence the interpretation of another (e.g., audio) based on their contextual relevance. This approach is particularly powerful because it supports fine-grained, temporally aligned interactions, capturing intricate dependencies across modalities that may evolve. These interactions make cross-modal attention especially well-suited for speech tasks, where audio and textual data are inherently synchronized and highly correlated. For instance, subtle prosodic cues in audio can enhance the interpretation of ambiguous textual content, while textual context can guide the model in focusing on relevant acoustic features. One typical implementation involves a cross-modal Transformer encoder that combines the key and value vectors of one modality with the query of another [275, 294]. This method, known as cross-attention [295], was first used in the VilBERT model [286]. In [95], the authors introduce a cross-modal attention Transformer that combines eye movement from video and audio features. This approach achieved an accuracy of 84.26% on their dataset, outperforming single modalitiy approaches.
@@
The late fusion strategy involves combining the results of each unimodal model after processing them separately [297]. This approach requires the development of separate models for each modality, allowing each to extract the most relevant information from its respective input. As a result, it provides a clearer understanding of each modality before their outputs are fused. The final fusion can be performed in various ways, for instance, by applying a weighted combination of modality-specific predictions. This combination is particularly beneficial in scenarios where one modality consistently outperforms the others. Alternatively, the fusion can be guided by the confidence scores of the models, assigning greater weight to predictions with higher confidence.
In [84], the authors propose using majority voting for classification tasks and a weighted average of the results for regression tasks on the ADReSS dataset. This methodology achieved an accuracy of 85.2%, 85.4% F1-Score, and a RMSE score of 4.65. Similarly, the work proposed in [81] also employs majority voting for their unimodal approaches. In this study, this fusion strategy yielded 86% accuracy on the same dataset, outperforming the unimodal approaches.
A joint or hybrid fusion strategy involves combining features obtained from each modality for later use in a final model, but only after being processed individually by each modality model [298]. These intermediate representations are then combined and passed to a final multimodal model for joint learning. This approach leverages the strengths of unimodal processing, allowing each modality-specific model to extract the most salient and task-relevant features from its input before any fusion occurs.
By combining these rich, modality-specific features at an intermediate level, the joint fusion strategy enables the final multimodal model to learn complex cross-modal relationships and interactions more effectively than early or late fusion alone. However, this strategy comes with increased computational cost and complexity. Since it relies on multiple components, two or more unimodal encoders followed by a multimodal model, it demands more memory and training resources than other fusion strategies. Despite this overhead, this combination is the most common approach in the selected works for this study.